OpenStudy (rebeccapink6398):

WILL MEDAL+FANNN Sharon is the youngest in her family. Her sister Anne is 3 years older than she is. Her brother Tom is 5 years older than Anne. The sum of their ages is 62. How old is Sharon?

OpenStudy (anonymous):

{s=a−3,s=t−5,a+s+t=62}

OpenStudy (rebeccapink6398):

?

OpenStudy (anonymous):

shanon is 18

OpenStudy (anonymous):

Shannon is 17. Not 18. I just did a guess and check. But 17+3 is 20. 20+5 =25. 17+20+25 = 62.

OpenStudy (rebeccapink6398):

thank you!!!! :D

OpenStudy (anonymous):

You could do it by making an equation. x = Shannon's age. x+3 = Anne's age. x+8 = Tom. x+x+3+x+8 =62. 3x+11 = 62. 3x=51. 51/3 = x. x=17.

OpenStudy (rebeccapink6398):

thats what someone eles said i didnt understand a work they said but you explained ti well than you so much! gosh home work is annoying XD

OpenStudy (anonymous):

Agreed! lol. Don't suppose you know how to find domain and range of exponential functions? :p

OpenStudy (rebeccapink6398):

haha i can barley figure out if i should round up or down XD A car company said that its top five models had the following fuel efficiency ratings: 26.3 mpg, 28.95 mpg, 33.6 mpg, 35 mpg, 38.24 mpg. What is the average mpg rating of these five models, recorded to the appropriate degree of precision? A. 32 mpg B. 32.0 mpg C. 32.4 mpg D. 32.42 mpg

OpenStudy (rebeccapink6398):

i say C. just a guess what do you say?

OpenStudy (anonymous):

Add the 4 numbers. Then divide by 4. Round to the nearest tenth. Like so...

OpenStudy (anonymous):

i think it a but not sure

OpenStudy (rebeccapink6398):

ugh. it was A jeee :(